List the four equations of linear motions (also referred to as the Kinematics equations). In your notations, consider “u” to be the initial velocity of the body, “v” the final velocity,” t” the time interval, “a” the acceleration and “s” the distance covered by the body .
Answer #1

=> Assuming u= initial velocity va final velocity ay f c > u t = time interval a = acceleration Set S = dis place ment/clistaalso, Velocity (instantaneous Olusplaamunt time v= ds » ds= vat 8 from Sdr. Sone Sensore fear [s1$ = [uts ere] ut dat? S = UtUsing , Sn= utn+ Latn² = unt 1 ano Sn= u(n-1) ++ a(n-1) Son – Some = (unt of an) – (urma) a ta crey) = u(n=(wy) + 4a(n? –(94Sn- Shus = u(1) + 5 a (n=n41) (n+n-1) ASn = u + zła (1) (2n-1) -♡ So, the 4 equations of motion are 1st > 2nd → 3rd → 4th → V